# Large Language Models Are Struggle to Cope with Unreasonability in Math Problems

### v5

@@ Abstract
Recent research have demonstrated LLMs’ impressive performance in math and reasoning. However, the capacity of LLMs to address math problems under unconventional conditions, such as internal inconsistencies and flawed assumptions, remains largely unexplored. In this paper, we propose a novel benchmark Unreasonable Math Problem (UMP) designed to assess LLMs’ ability to recognize and respond to unreasonability in math problem. The benchmark consists of a carefully curated collection of unreasonable math questions across diverse types. Based on extensive experiments covering 19 LLMs, we observe that even state-of-the-art models such as GPT-4o achieve only limited performance of 0.6 in UMP, while reasoning models such as DeepSeek-R1 are prone to overthinking and unstable. We further explore strategies for improving the recognition of unreasonable inputs, shedding light on both the possibility and limitations of LLMs in this challenging setting.
Large Language Models Are Struggle to Cope with Unreasonability
in Math Problems
@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable progress in math and reasoning, particularly with the emergence of reasoning-enhanced models like O1 and R1 DeepSeek-AI (2025), demonstrate increasingly structured and deliberate reasoning capabilities. As a result, current LLMs have surpassed earlier baselines and achieved over 90% accuracy on math benchmarks like GSM8K Cobbe et al. (2021a) and MATH Hendrycks et al. (2021).
However, despite this impressive performance on challenging math problems, we find that these models often struggle to recognizing when a math question is unreasonable. Even when the irrationality, such as a logical contradiction or a violation of basic constraints, is immediately obvious to humans. LLMs frequently follow the misleading question setup, producing confident yet nonsensical answers, or becoming trapped in cycles of excessive reasoning that fail to converge. This counter-intuitive behavior raises serious concerns about the reliability in their practical applications such as automated tutoring Kasneci et al. (2023), early education Zhang et al. (2024b), or open-domain problem solving Lin and Chen (2023).
To enable a comprehensive analysis of how LLMs behave when confronted with mathematically unreasonable inputs, we introduce the Unreasonable Math Problems (UMP) benchmark. We take questions from GSM8K Cobbe et al. (2021a), MATH Hendrycks et al. (2021), AIME24 MAA (2024), and AMC23 to generated unreasonable variants. Specifically, we utilize GPT-4o to apply minimal edits that introduce specific forms of unreasonability while preserving the original questions structure. These edits are guided by a set of rule-based transformation templates. All generated questions are verified to ensure that they are genuinely unreasonable in a mathematical or logical sense, and that they remain similar to the original questions in length, topic, and surface form. This careful construction allows us to attribute differences in model behavior specifically to the presence of unreasonableness, rather than to confounding changes in difficulty.
Based on UMP benchmark, we conducted extensive experiments on various models spanning three categories: general-purpose models (e.g., GPT-4oOpe (2023)), reasoning-enhanced models (e.g., DeepSeek-R1DeepSeek-AI (2025)), and math-specialized models (e.g., Qwen-MathYang et al. (2024a)). In addition, we analyze LLMs’ failure patterns according to tokens repetition, reflection frequency, and token entropy. We find: 1) general-purpose models often proceed confidently without recognizing the unreasonableness of the question; 2) reasoning-enhanced models tend to overthink and fall into excessive self-correction; 3) math-specialized models may fail to initiate reasoning when confronted with unreasonable premises. We further explored several mitigation strategies but found that none could robustly resolve these failure modes without introducing trade-offs, such as decreased performance on standard inputs.
@@
We propose the Unreasonable Math Problems (UMP) benchmark by minimally editing standard math datasets to create diverse and comprehensive unreasonable problems.
We conduct comprehensive experiments and multi-dimensional analyses to examine how different types of LLMs behave when presented with unreasonable questions.
We identify distinct failure modes and reveal that models with strong performance on well-posed tasks often behave unstably or fail to recognize unreasonableness when assumptions are flawed.
We explore prompting and fine-tuning as strategies to investigate the potential of LLMs to detect unreasonableness.
@@ 2 Unreasonable Math Problems(UMP) Benchmark
While large language models have demonstrated strong performance on standard mathematical reasoning benchmarks, they often produce inaccurate or confusing responses when presented with mathematically unreasonable problems, questions that contain flawed assumptions, undefined variables, or logical inconsistencies (as show in Figure 1). To systematically evaluate model behavior under such conditions, we construct the Unreasonable Math Problems (UMP) benchmark, which consists of 1000+ unreasonable math problems, each paired with its corresponding original version, focusing on assessing LLMs’ ability to detect and respond to irrational inputs.
2.1 Data Category
@@
We identify five prevalent types of mathematical unreasonableness commonly found in LLM failure cases: (1) undefined variables, (2) illogical scenarios, (3) incorrect assumptions, (4) misinterpretation of units, and (5) inconsistent conditions. Each instance in our benchmark is represented as a quintuple , consisting of the original question and its answer , the unreasonable variant with its assigned type , and an explanation describing why is unreasonable. Detailed definitions and examples for each category are provided in Appendix A.
2.1.1 Data Generation
@@
Our data generation process is inspired by MetaMath (Yu et al., 2023b), which leverages LLMs to produce problem variants under controlled transformations. As illustrated in Figure 2, we begin with test set questions drawn from four widely-used math benchmarks: GSM8K Cobbe et al. (2021a), a collection of grade-school level problems; MATH Hendrycks et al. (2021), which covers formal secondary school mathematics; and AIME24MAA (2024) and AMC23, both of which contain high level competition problems with symbolic or abstract formats. We manually construct a set of transformation rules corresponding to five types of mathematical unreasonableness and use them to guide an LLM in generating unreasonable variants for each original question, along with natural language explanations of why the modified version is irrational. To ensure the unreasonable variant remains close in surface form to the original, we compute cosine similarity between the sentence embeddings of the original and generated questions using SimCSE (Gao et al., 2021), a contrastively trained BERT-based modelDevlin et al. (2019). Only variants whose similarity exceeds a predefined threshold are retained for human verification.
2.1.2 Data Checking
@@
We perform manual validation with the assistance of three annotators, each with at least undergraduate-level mathematical training. For each example, annotators are provided with the original problem, its unreasonable variant, and a corresponding explanation, and are asked to verify whether the variant introduces genuine mathematical unreasonableness while preserving surface similarity. Only examples that meet both criteria are retained in the final benchmark. Our transformation rules are deliberately designed to introduce flaws that are logically problematic. As such, we filter out cases that either fail to introduce meaningful irrationality or make the unreasonableness too explicit, for example, variants that contain phrases like “-3 people” or “2.5 people,” which violate commonsense constraints in an obvious way. In these cases, we revise the questions by modifying entities, adjusting numerical values, or refining the phrasing to ensure that the unreasonable elements remain logically embedded but superficially plausible.
2.1.3 Data Composition
@@
Table 1 summarizes the distribution of unreasonableness types across different source datasets in the UMP benchmark. A large proportion of examples fall into the categories of Incorrect Assumptions (IA) and Inconsistent Conditions (IC). This distribution emerges from the model generation and human filtering process. In particular, these two types of flaws tended to produce more plausible and contextually coherent questions, making them more likely to be retained during human recheck. In comparison, AMC23 and AIME24 contain relatively few test questions (with AIME24 having only 30), which naturally limits the number of examples that can be derived from these datasets. Moreover, problems in these sources are often abstract and symbolic, making them less amenable to realistic and controlled perturbation. Rather than artificially forcing unreasonable variants in such cases, we adopt a conservative strategy and include only those examples where natural modifications can be applied without compromising the original structure. This selective inclusion ensures that the final benchmark maintains both diversity and quality across a range of mathematical domains.
@@
3 Experiments
3.1 Experimental Setup
We evaluate a diverse set of LLMs spanning three major categories. General-purpose models are primarily trained for instruction following and open-domain tasks; this group includes DeepSeek-V3 DeepSeek-AI (2024), GPT-4o Ope (2023), Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct Qwen Team (2024), and LLaMA3.1-8B-Instruct AI (2024). Math-specialized models are fine-tuned on mathematical corpora and optimized for numerical reasoning; we include Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct Yang et al. (2024a), and DeepSeek-Math-7B-Instruct Shao et al. (2024). Reasoning-enhanced models are designed to support complex multi-step reasoning and include DeepSeek-R1 and its distilled variants (e.g., DeepSeek-R1-Distill-Qwen-7B) DeepSeek-AI (2025), as well as QwQ-32B-Preview Team (2025) and Grok-3-Reasoning. All models are evaluated on both the original problems and their unreasonable variants.111Details of model versions and inference configurations are provided in Appendix C.
@@ 3.2 Evaluation Metrics
Evaluating model performance on unreasonable questions poses a fundamentally different challenge compared to standard accuracy-based evaluation. In conventional settings, the goal is to assess whether the model arrives at the correct answer. However, for unreasonable questions, the objective is to evaluate whether the model can recognize the irrationality embedded in the problem and respond appropriately by rejecting, questioning, or flagging the input as flawed. To address this, we introduce two complementary metrics:
@@
Relative score conditions on the model’s ability to correctly solve the original version of a question. It is defined as the proportion of corresponding unreasonable variants that are correctly recognized as unreasonable, among those for which the model answers the original (reasonable) version correctly. This design controls for a factor that in some cases, the irrationality in a problem only becomes apparent during intermediate reasoning steps. A model that lacks the necessary mathematical competence may never reach the point where the irrationality is revealed. By restricting the evaluation to questions the model can already solve in their original form, relative score isolates the model’s ability to detect unreasonableness from its general problem-solving skill.
@@
| Qwen2.5-Math-1.5B-Instruct | 0.376 | |||||||||||
@@
| Qwen2.5-Math-7B-Instruct | ||||||||||||
@@ 3.2.1 Absolute Score
To evaluate how well a model identifies the unreasonableness of a given question, we adopt the LLM-as-a-judge paradigm Zheng et al. (2023), using DeepSeek-V3 to assess the quality of generated responses. Instead of a binary correct/incorrect judgment, we use a three-level rubric to better capture the nuances of model behavior:
A: Correctly identifies the main source of unreasonableness, provides a coherent explanation, and proposes a valid correction without introducing new errors.
B: Partially recognizes the unreasonableness or provides vague justification, but the explanation contains gaps, circular logic, or minor inconsistencies.
C: Fails to recognize the core flaw in the problem, misinterprets the reasoning, or introduces new contradictions in the explanation or correction.
@@ 3.2.2 Relative Score
While the absolute score provides a general measure of robustness, it does not account for the model’s underlying mathematical competence. In many cases, the flaw in an unreasonable question only becomes apparent during intermediate reasoning steps. If a model lacks the skill to solve the original problem, it may never reach the point where the flaw is revealed.
To control for this confound, we introduce the relative score, which measures a model’s ability to detect unreasonableness on the subset of questions it has already solved correctly in their original form. Formally, following Yang et al. (2024b), we define the relative score as:
@@ 3.3 Experimental Results
General-purpose models.
Models such as GPT-4o, Claude-3.5-Sonnet, and DeepSeek-v3 perform well on well-posed problems (e.g., GPT-4o: 0.863 Acc) and show moderate ability to detect unreasonableness (0.603 Abs). However, smaller models like Qwen2.5-3B and LLaMA-3.1-8B struggle with behavioral metrics, indicating difficulty in recognizing flawed logic. This suggests that robustness to irrationality is not a natural outcome of instruction tuning and may require additional improvement.
Reasoning-enhanced models.
These models achieve consistently high scores, with DeepSeek-R1 and grok-3-reasoning reaching 0.813 Abs and 0.903 Rel, respectively. Their structured reasoning and verification steps appear effective for detecting logical flaws, though often at the cost of longer, more verbose outputs (see Section 4.2).
Math-specialized models.
Despite strong accuracy on original problems (e.g., Qwen2.5-Math-7B: 0.894 Acc), this group generally underperforms on unreasonable problems(score mostly below 0.30), revealing a gap between reasoning proficiency and irrationality detection. MetaMath-Mistral-7B exemplifies this limitation, with only 0.139 on Rel.
Summary.
We observe a trade-off between task accuracy and robustness to unreasonableness. Reasoning-enhanced models excel at flaw detection, while math-specialized models, despite high precision on well-posed problems, often fail to recognize inconsistencies in adversarial variants.
@@ 4.2 Reasoning Models Tend to Overthink on Unreasonable Problems
A prominent behavioral failure we observe is overthinking, where the model enters repeated cycles of reflection and revision without making meaningful progress Sui et al. (2025); Cuadron et al. (2025). We quantify overthinking by counting occurrences of reflective phrases such as “rethink” or “I misunderstood,” which indicate mid-generation self-correction (see Appendix D for keyword details). Although we evaluate reflection frequency across all model types, only reasoning-enhanced models exhibit consistent and substantial increases on unreasonable problems.
@@ 4.3 Lexical Collapse and Redundancy in Response
As shown in Figure 4, most models exhibit a decline in entropy on unreasonable questions. This drop is especially pronounced in reasoning-enhanced models such as R1-Distill-Qwen-32B, where entropy decreases by 0.1–0.15 on average. These collapsed outputs often stem from repeated, ineffective reasoning loops, which reduce informativeness and mask the model’s failure to recognize flawed assumptions. Examples of such behavior are presented in Appendix I.1.
@@
Figure 4 shows that almost all models exhibit increased 10-gram repetition under unreasonable inputs, with the most substantial jumps observed in reasoning-enhanced models. In some cases, repetition increases by a factor of 5 to 10 compared to original problems, reinforcing the pattern of collapsed outputs. Interestingly, Qwen2.5-Math-7B-Instruct, a math-specialized model with strong performance on well-posed questions, also shows a marked rise in repetition, suggesting that semantic instability is not limited to reasoning-heavy models. Appendix I.2 provides detailed case studies of this failure pattern.
5 Can LLMs Detect Irrationality?
Through the experiments above, we observe that most models perform poorly at recognizing unreasonableness in mathematical problems. To explore possible ways to improve this behavior, we begin with a simple probing experiment to test whether models possess the basic ability to judge flawed inputs. Based on this, we further investigate two strategies, prompt-based intervention and supervised fine-tuning—to enhance the model’s capacity to detect irrationality.

### v6

# Large Language Models Struggle with Unreasonability in Math Problems

@@ Abstract
Large Language Models (LLMs) have shown remarkable success on a wide range of math and reasoning benchmarks. However, we observe that they often struggle when faced with unreasonable math problems. Instead of recognizing these issues, models frequently proceed as if the problem is well-posed, producing incorrect answers or falling into overthinking and verbose self-correction. To systematically investigate this overlooked vulnerability, we propose the Unreasonable Math Problems (UMP) benchmark, designed to evaluate LLMs’ ability to detect and respond to unreasonable math problem statements. Based on extensive experiments covering 19 LLMs, we find that even state-of-the-art general models like GPT-4o achieve only a score of 0.6 on UMP. While reasoning models such as DeepSeek-R1 demonstrate a higher sensitivity to unreasonable inputs, this often comes at the cost of generating overly long and meaningless responses that fail to converge. We further explore prompting and fine-tuning methods, which offer partial improvements but also introduce trade-offs, shedding light on both the potential and limitations of LLMs in this challenging setting.
Large Language Models Struggle with Unreasonability in Math Problems
@@ 1 Introduction
Large language models (LLMs) have recently shown impressive performance on advanced mathematical reasoning tasks, especially on benchmarks like MATH (Hendrycks et al., 2021) and AIME24MAA (2024). However, we find that these models often fail to detect logical flaws or unreasonable assumptions in math problems, treating them as if they were well-posed. Instead of flagging such issues, they tend to generate confident yet nonsensical answers, or fall into endless reasoning loops without reaching a valid conclusion. This counter-intuitive behavior raises serious concerns about their reliability in real-world applications such as automated tutoring Kasneci et al. (2023), early education Zhang et al. (2024b), and open-domain problem solving Lin and Chen (2023), where misleading answers to unreasonable questions can undermine trust and lead to negative outcomes.
To enable a comprehensive analysis of how LLMs behave when confronted with mathematically unreasonable inputs, we introduce the Unreasonable Math Problems (UMP) benchmark. We construct UMP by minimally editing questions from existing math datasets, including MATH Hendrycks et al. (2021), AIME24 MAA (2024), AMC23, and GSM8K Cobbe et al. (2021a) to create unreasonable variants that contain logical inconsistencies, missing assumptions, or ill-defined objectives. These edits are guided by rule-based transformation templates and executed using DeepSeek-R1, while preserving the original problem’s structure, topic, and surface form. All generated questions are manually verified to ensure they are mathematically flawed yet still natural.
To evaluate model behavior, we present models with both the original and its corresponding unreasonable version. This paired setup enables us to directly attribute behavioral differences to the presence of unreasonableness, isolating it from other factors such as question length or difficulty.
Based on the UMP benchmark, we conduct extensive experiments on various models spanning three categories: general-purpose models (e.g., GPT-4oOpenAI (2024)), reasoning models (e.g., DeepSeek-R1DeepSeek-AI (2025)), and math-specialized models (e.g., Qwen-MathYang et al. (2024a)). In addition, we analyze LLMs’ failure patterns according to tokens repetition, reflection frequency, and token entropy. We find: 1) general-purpose models often proceed confidently without recognizing the unreasonableness of the question; 2) reasoning models tend to overthink and fall into excessive self-correction; 3) math-specialized models may fail to initiate reasoning when confronted with unreasonable premises. We further explore several mitigation strategies but find that none could robustly resolve these failure modes without introducing trade-offs, such as decreased performance on standard inputs, highlighting the need for future research into more principled and generalizable solutions
@@
We propose the Unreasonable Math Problems (UMP) benchmark to more accurately and comprehensively evaluate how LLMs respond to mathematically unreasonable problem statements.
We find that even high-performing models often fail to detect unreasonableness, or produce overconfident and overly verbose responses to unreasonable questions.
We show that simple prompting or fine-tuning can partially mitigate these issues, but often introduce trade-offs such as decreased accuracy on well-posed problems, posing new challenges for future research.
@@ 2 Unreasonable Math Problems(UMP) Benchmark
While large language models have demonstrated strong performance on standard mathematical reasoning benchmarks, they often produce inaccurate or confusing responses when presented with mathematically unreasonable problems, questions that contain flawed assumptions, undefined variables, or logical inconsistencies (as shown in Figure 1). To systematically evaluate model behavior under such conditions, we construct the Unreasonable Math Problems (UMP) benchmark, which consists of 1000+ unreasonable math problems, each paired with its corresponding original version, focusing on assessing LLMs’ ability to detect and respond to irrational inputs.
2.1 Types of Unreasonableness
@@
We identify five prevalent types of mathematical unreasonableness commonly found in LLM failure cases: (1) undefined variables, (2) illogical scenarios, (3) incorrect assumptions, (4) misinterpretation of units, and (5) inconsistent conditions. Each instance in our benchmark is represented as , consisting of the original question and its answer , the unreasonable variant with its assigned type , and an explanation describing why is unreasonable. Detailed definitions and examples for each category are provided in Appendix A.
2.2 LLM-Guided Construction of Unreasonable Variants
@@
Our data construction process is inspired by MetaMath (Yu et al., 2023b), which leverages LLMs to produce problem variants under controlled transformations. As illustrated in Figure 2, we begin with test set questions drawn from four widely-used math benchmarks: GSM8K Cobbe et al. (2021a), a collection of grade-school level problems; MATH Hendrycks et al. (2021), which covers formal secondary school mathematics; and AIME24MAA (2024) and AMC23, both of which contain high level competition problems with symbolic or abstract formats. We manually construct a set of transformation rules corresponding to five types of mathematical unreasonableness and use them to guide an LLM in generating unreasonable variants for each original question, along with natural language explanations of why the modified version is irrational. To ensure the unreasonable variant remains close in surface form to the original, we compute cosine similarity between the sentence embeddings of the original and generated questions using SimCSE (Gao et al., 2021), a contrastively trained BERT-based modelDevlin et al. (2019). Only variants whose similarity exceeds a predefined threshold are retained for human verification.
2.3 Validation Checking
@@
We manually verify each example to ensure that the unreasonable variant introduces genuine mathematical unreasonableness while maintaining surface similarity with the original problem. Only examples satisfying both criteria are retained. Our transformation rules are carefully designed to embed logical flaws. We discard cases that either lack meaningful irrationality or make the flaw overly explicit (see the annotation protocol in Appendix H). In such cases, we revise the examples by adjusting entities, numerical values, or phrasing to ensure the unreasonableness remains logically subtle yet plausible.
2.4 Benchmark Composition
@@
Table 1 summarizes the distribution of unreasonableness types across different source datasets in the UMP benchmark. A large proportion of examples fall into the categories of Incorrect Assumptions (IA) and Inconsistent Conditions (IC). This distribution emerges from the model generation and human filtering process. In particular, these two types of flaws tend to produce more plausible and contextually coherent questions, making them more likely to be retained during human recheck. Compared to GSM8K and MATH, the AMC23 and AIME24 datasets contain far fewer test questions (e.g., AIME24 has only 30), which inherently limits the number of examples we can derive from them. In addition, these problems are often highly abstract and symbolic, making it difficult to apply natural, controlled perturbations without compromising their integrity. As a result, we only include modified examples when the unreasonableness can be introduced in a plausible manner. This selective inclusion ensures that the final benchmark remains both diverse and faithful to the original problem distributions.
@@
3 Evaluating LLMs on Unreasonable Math Problems
3.1 Evaluation Setup
We evaluate a diverse set of LLMs spanning three major categories. General-purpose models are primarily trained for instruction following and open-domain tasks; this group includes DeepSeek-V3 DeepSeek-AI (2024), GPT-4o Ope (2023), Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct Qwen Team (2024), and LLaMA3.1-8B-Instruct AI (2024). Math-specialized models are fine-tuned on mathematical corpora and optimized for numerical reasoning; we include Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct Yang et al. (2024a), and DeepSeek-Math-7B-Instruct Shao et al. (2024). Reasoning-enhanced models are designed to support complex multi-step reasoning and include DeepSeek-R1 and its distilled variants (e.g., DeepSeek-R1-Distill-Qwen-7B) (DeepSeek-AI, 2025), as well as QwQ-32B-Preview (Team, 2025) and Grok-3-Reasoning (xAI, 2024). All models are evaluated on both the original problems and their unreasonable variants.111Details of model versions and inference configurations are provided in Appendix D.
@@ 3.2 Evaluation Metrics
Evaluating model performance on unreasonable questions poses a fundamentally different challenge compared to standard accuracy-based evaluation. In conventional settings, the goal is to assess whether the model arrives at the correct answer. However, for unreasonable questions, the objective is to evaluate whether the model can recognize the unreasonability embedded in the problem and respond appropriately by rejecting, questioning, or flagging the input as flawed. To address this, we introduce two complementary metrics:
@@
Relative score conditions on the model’s ability to correctly solve the original version of a question. It is defined as the proportion of corresponding unreasonable variants that are correctly recognized as unreasonable, among those for which the model answers the original (reasonable) version correctly. This design accounts for cases where the flaw only becomes apparent during intermediate reasoning steps. A model that lacks the necessary mathematical competence may never reach the point where the unreasonability is revealed. By restricting the evaluation to questions the model can already solve in their original form, relative score isolates the model’s ability to detect unreasonableness from its general problem-solving skill.
@@
| Qwen2.5-Math-1.5B-Instruct | 0.356 | 0.416 | 0.376 | |||||||||
@@
| Qwen2.5-Math-7B-Instruct | 0.458 | |||||||||||
@@ 3.2.1 Absolute Score
Following the LLM-as-a-judge framework Zheng et al. (2023), we use DeepSeek-V3 to label each answer as A (correctly identifies the main source of unreasonableness, explains it coherently, and proposes a valid fix without new errors), B (partially detects or vaguely justifies the flaw, with gaps, circular logic, or minor inconsistencies), or C (fails to spot the core flaw, misreads the reasoning, or introduces new contradictions).
@@ 3.2.2 Relative Score
If a model lacks the skill to solve the original problem, it may never reach the point where the flaw is revealed. To control this factor, we introduce the relative score, which measures a model’s ability to detect unreasonableness on the subset of questions it has already solved correctly in their original form. Formally, following Yang et al. (2024b), we define the relative score as:
@@ 3.3 Experimental Results
We categorize evaluated models into three groups: general-purpose, reasoning-enhanced, and math-specialized. General models such as GPT-4o, Claude-3.5-Sonnet, and DeepSeek-v3 achieve strong accuracy on well-posed problems (e.g., GPT-4o: 0.863 Acc) and moderate robustness to unreasonableness (0.603 Abs), while smaller models like Qwen2.5-3B and LLaMA-3.1-8B struggle across behavioral metrics. Reasoning models (e.g., DeepSeek-R1, grok-3-reasoning) perform best on UMP, reaching 0.813 Abs and 0.903 Rel, benefiting from structured verification steps, but often generating overly verbose responses (see Section 4.1). In contrast, math-specialized models (e.g., Qwen2.5-Math-7B, MetaMath-Mistral-7B) excel in original task accuracy (up to 0.894 Acc) but perform poorly on UMP (mostly below 0.30), indicating a gap between precise reasoning and unreasonability detection.
These results highlight a trade-off between accuracy on well-posed problems and sensitivity to unreasonableness, with reasoning models offering better detection at the cost of increased verbosity.
@@ 4.1 Reasoning Models Tend to Overthink on Unreasonable Problems
A prominent behavioral failure we observe is overthinking, where the model enters repeated cycles of reflection and revision without making meaningful progress Sui et al. (2025); Cuadron et al. (2025). We quantify overthinking by counting occurrences of reflective phrases such as “rethink” or “I misunderstood,” which indicate mid-generation self-correction (see Appendix F for keyword details). Although we evaluate reflection frequency across all model types, only reasoning-enhanced models exhibit consistent and substantial increases on unreasonable problems.
@@ 4.2 Lexical Collapse and Redundancy in Response
As shown in Figure 4, most models exhibit a decline in entropy on unreasonable questions. This drop is especially pronounced in reasoning-enhanced models such as R1-Distill-Qwen-32B, where entropy decreases by 0.1–0.15 on average. These semantic collapse often stems from repeated, ineffective reasoning loops, which reduce informativeness and mask the model’s failure to recognize flawed assumptions. Examples of such behavior are presented in Appendix L.1.
@@
Figure 4 shows that almost all models exhibit increased 10-gram repetition under unreasonable inputs, with the most substantial jumps observed in reasoning-enhanced models. In some cases, repetition increases by a factor of 5 to 10 compared to original problems, reinforcing the pattern of collapsed outputs. Interestingly, Qwen2.5-Math-7B-Instruct, a math-specialized model with strong performance on well-posed questions, also shows a marked rise in repetition, suggesting that semantic instability is not limited to reasoning-heavy models. Appendix L.2 provides detailed case studies of this failure pattern.
4.3 Summary of Findings
Our behavioral analysis uncovers several key patterns. First, general-purpose models often proceed with unwarranted confidence, failing to recognize flawed assumptions—a phenomenon we term unconscious of unreasonableness. Second, reasoning-enhanced models are prone to overthinking, repeatedly revising their reasoning in response to irrational inputs, which leads to verbosity and incoherence. Third, both reasoning and math-specialized models exhibit semantic collapse, characterized by increased token repetition and reduced entropy under unreasonable conditions. Finally, despite strong accuracy on well-posed tasks, math-specialized models often fail to detect subtle logical flaws, highlighting a disconnect between mathematical proficiency and robustness to flawed inputs.
5 Can LLMs Detect unreasonability?
Through the experiments above, we observe that most models perform poorly at recognizing unreasonableness in mathematical problems. To explore possible ways to improve this behavior, we begin with a simple probing experiment to test whether models possess the basic ability to judge flawed inputs. Based on this, we further investigate two strategies, prompt-based intervention and supervised fine-tuning to enhance the model’s capacity to detect unreasonability.